OpenStudy (anonymous):

Hey there I need help with solving for x equation. Y= -2 X+2y=7/4 I keep getting 23/4 but it is not the correct answer

OpenStudy (abb0t):

x+2(-2)= \(\sf \color{}{\frac{7}{4}}\) \(\Rightarrow\) x-4= \(\sf \color{}{\frac{7}{4}}\)

OpenStudy (anonymous):

Nope still wrong

OpenStudy (phi):

Are you sure about the equation? if you really have y= -2 and \[ x+2y = \frac{7}{4} \] then x is 23/4

OpenStudy (abb0t):

I am never wrong. Just let that sink in.

OpenStudy (anonymous):

But they are saying the answer is 15/2

OpenStudy (abb0t):

Then they are wrong.

OpenStudy (phi):

Ask your teacher about this one.

OpenStudy (phi):

The question probably has a typo in it.

OpenStudy (anonymous):

Probably I'm gonna check it again sighhh

OpenStudy (phi):

Take a screen capture of the question and show your teacher.

OpenStudy (anonymous):

yeah this question is awkward they did make a typo in for the question y should be -4 not - 2 its not possible

OpenStudy (anonymous):

wait I was right it is 23/4 i just miscalculated when i wanted to substitute for x ugh